answer for x^2+2x+5=0
Answer by stanbon(75887)

x^2+2x+5=0
Use the quadratic formula to get:
x=[-2+-sqrt(4-4*5)]/2
x=[-2+-sqrt(-16)]/2
x=[-2+-4i]/2
x=-1+2i or x=-1-2i
---------
Cheers,
Stan H.